OpenStudy (anonymous):

Solve for variable x: 5=12/x (5 equals 12 over x) A. 5/12 B. 4.12 C. 12/5 D. 2.75

OpenStudy (zephyr141):

multiply both sides by x then divide both sides by 5.

OpenStudy (zephyr141):

\[5=\frac{ 12 }{ x }\] \[5*x=\frac{ 12 }{ x }*x\] here x divides out on the right side of the equation. \[\frac{ 5x }{ 5 }=\frac{ 12 }{ 5 }\] here it's the same with five on the left side \[x=\frac{ 12 }{ 5 }\]finally you're left with x

OpenStudy (triciaal):

the LCD of 1 and x = x so multiply each side by x to get 1 x divide by the coefficient if you have 5x divide by 5 to get x if you had 10 a = 20 you would divide by 10 to get a = 2

OpenStudy (anonymous):

thank you both. appreciate it